# arXiv:2603.00989v2[cs.SE] 03 Mar 2026

### v2

@@ IV-C1 Evaluation Metrics
Table XI summarizes the evaluation metrics reported across the selected studies. It can be observed that most studies focus on energy consumption and performance metrics, particularly latency and runtime. Fewer studies evaluate functional correctness and memory usage, and only a small number consider code quality.
@@
Energy Energy consumption [96], [110], [22], [27], [60], [25], [105], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10]
@@
Performance Latency/Runtime [96], [110], [22], [27], [60], [25], [105], [130], [129], [64], [101], [11], [97], [116], [12]
@@
TABLE XI: Categorization of evaluation metrics reported in the selected studies.
IV-C2 Custom Evaluation Metrics
In addition to standard evaluation metrics, a small number of studies introduce custom sustainability-oriented metrics. These metrics are designed by the authors to capture specific aspects of energy efficiency, carbon impact, or sustainability that are not covered by conventional measures. Table XII presents the studies that propose such custom evaluation metrics and lists the corresponding metrics introduced in each work. Studies that do not propose any custom evaluation metrics are omitted from this table.
Study Custom Evaluation Metrics
[127] RuntimeRatio, EnergyRatio
[33] Energy-reduction@k
[129] Green Capacity (GC)
TABLE XII: Studies proposing custom sustainability-oriented evaluation metrics.
RuntimeRatio and EnergyRatio compare the runtime and energy consumption of model-generated code against canonical human-written solutions, providing a relative efficiency measure. Energy-reduction@k estimates the expected energy savings when generating multiple candidate solutions and selecting the most energy-efficient valid one, while accounting for both performance improvement and correctness probability. Similarly, Green Capacity (GC) measures the overall sustainability improvement of optimized code across multiple performance dimensions, filtering out invalid or degraded solutions. Embodied Energy and Embodied Carbon capture the environmental impact associated with code generation or development infrastructure, whereas Operational Energy and Operational Carbon measure emissions during actual code execution. Carbon Intensity (CI) quantifies the amount of carbon emissions per unit of electricity consumed.
IV-C3 Libraries / Tools for Measurement
@@ IV-C4 Hardware Configuration
Evaluation Hardware Category Representative Configurations
On-premise (local machine/server) [110], [127], [22], [27], [60], [129], [64], [116], [10]
@@ IV-C5 Experimental Configuration
IV-C6 Open-Source Availability
@@
Category Count
Open-source available [96], [110], [33], [22], [27], [60], [25], [105], [64], [97], [116], [29], [10]
@@
IV-C7 Benchmark Comparison and Usage
We also categorize the evaluation data based on the type of benchmark or dataset used and the programming language considered. Table XVII summarizes the benchmarks and datasets used in the selected studies, along with programming language information when explicitly reported. Programming languages are reported only when explicitly specified by the authors or when the language choice is central to the evaluation; for standardized benchmarks, the programming language is implicit and therefore omitted. It can be observed that most studies primarily focus on Python, reflecting its widespread use in code generation research. In addition, not all studies rely on standardized benchmarks, as several works evaluate models using diverse or custom data sources, indicating variability in evaluation practices across the literature.
@@ IV-C7 Benchmark Comparison and Usage
IV-D Model Adaptation
IV-D1 Fine-Tuning
We also categorize the studies based on the use of fine-tuning, that is, whether a study adapts the model parameters to generate more sustainable code. Surprisingly, only one study [60] employs fine-tuning for this purpose. This study proposes a fine-tuning approach for code generation models using a weighted aggregated loss [77], which enables models to perform dynamic early exits at different layers without requiring additional language model heads. In addition, it introduces GREEN-CODE, a reinforcement learning–based [71] framework that improves energy efficiency in code generation by dynamically selecting early exit points during inference.
IV-D2 Prompt Engineering
Table XVIII summarizes whether the selected studies use prompt engineering techniques. A study is marked as “Yes” if it explicitly applies methods such as zero-shot, few-shot, chain-of-thought, iterative prompting, or prompt variations to influence code generation. Studies that use default model settings without modifying prompts are marked as “No.” It can be observed that several studies rely on prompt-based strategies, while others evaluate models without applying specific prompt engineering techniques.
@@ IV-D2 Prompt Engineering
No [127], [60], [25], [105], [130], [129], [64], [11], [116]
TABLE XVIII: Use of prompt engineering techniques in the selected studies.
V Future Research Opportunities
@@
V-A Expanding Sustainability Beyond Code Generation
@@
V-B Underexplored Application Domains
@@
Future research should explore sustainability-aware code generation across a broader range of application domains. In particular, mobile and client-side applications, low-level and systems software, web and cloud infrastructure, and IoT and embedded systems are not adequately represented in the selected studies. These domains are especially important from a sustainability perspective, as energy efficiency directly affects battery life, system performance, scalability, and overall environmental impact. For example, embedded and IoT systems operate under strict energy constraints, low-level software can significantly influence hardware efficiency, and cloud-based services contribute substantially to large-scale energy consumption. The limited attention to these areas highlights the need for future studies to evaluate LLM-generated code in more realistic, energy-sensitive, and deployment-oriented environments.
V-C Expanding Sustainability Evaluation Across Programming Languages
@@
This direction is particularly important from a sustainability perspective, as energy efficiency can vary significantly across programming languages and execution environments. Different languages introduce different abstractions, memory management mechanisms, and runtime behaviors, all of which influence energy consumption. Moreover, language choice is often tightly coupled with the target platform, such as mobile devices, cloud infrastructure, edge systems, blockchain networks, or networked distributed systems, which further shapes optimization opportunities, runtime characteristics, and energy profiles.
@@
V-D Toward Sustainability-Aware LLM Optimization
Another important direction for future research is the development of sustainability-aware fine-tuning and model adaptation techniques. Our review shows that only one study [60] applies fine-tuning or reinforcement learning–based adaptation with an explicit focus on improving energy efficiency. This suggests that most existing work measures sustainability at the output level without modifying the underlying model to generate more energy-efficient code.
Future research should explore energy-aware fine-tuning strategies specifically tailored for sustainable code generation and optimization tasks. Instead of only measuring the energy consumption of generated code after inference, models should be trained to directly produce energy-efficient implementations. For example, training objectives can be extended to penalize inefficient code patterns, high-complexity logic, or unnecessary computational operations, encouraging the generation of optimized and resource-efficient solutions. Green reinforcement learning approaches may also be applied, where models receive reward signals based on improvements in runtime efficiency, reduced energy consumption, or lower carbon footprint of the generated code. Moreover, parameter-efficient adaptation methods can further reduce training and deployment cost, making sustainability-aware code generation more practical. By embedding sustainability objectives directly into model training and inference processes, future research can enable LLMs to proactively generate optimized, energy-efficient code rather than relying solely on post-hoc energy evaluation.
V-E Limited Focus on Small Models
@@
V-F Need for Sustainability-Oriented Benchmarks
Another important observation from this study is that none of the selected works introduce a benchmark specifically designed to evaluate sustainability. Instead, all studies use existing general-purpose benchmarks and add sustainability-related metrics, such as energy consumption or carbon footprint, as additional evaluation measures. This indicates a clear research gap in the field. Future research should focus on developing standardized benchmarks dedicated to sustainability-aware code generation. These benchmarks should include tasks and evaluation methods that measure energy efficiency, carbon impact, resource usage, and the ability of models to produce energy-efficient solutions. Creating such benchmarks would improve comparability across studies and support more consistent and rigorous evaluation of sustainable code generation approaches.
V-G Energy Measurement and Hardware Evaluation Gaps

### v3

@@ V-C1 Evaluation Metrics
Study Custom Evaluation Metrics
[122] RuntimeRatio, EnergyRatio
[32] Energy-reduction@k
[124] Green Capacity (GC)
TABLE XI: Studies proposing custom sustainability-oriented evaluation metrics.
@@
Energy Energy consumption [92], [105], [21], [26], [24], [101], [125], [124], [62], [97], [10], [93], [111], [28], [11], [9]
@@
Performance Latency/Runtime [92], [105], [21], [26], [24], [101], [125], [124], [62], [97], [10], [93], [111], [11]
@@
TABLE XII: Categorization of evaluation metrics reported in the selected studies.
Table XII summarizes the evaluation metrics reported across the selected studies. It can be observed that most studies focus on energy consumption and performance metrics, particularly latency and runtime. Fewer studies evaluate functional correctness and memory usage, and only a small number consider code quality.
V-C2 Custom Evaluation Metrics
In addition to standard evaluation metrics, a small number of studies introduce custom sustainability-oriented metrics. These metrics are designed by the authors to capture specific aspects of energy efficiency, carbon impact, or sustainability that are not covered by conventional measures. Table XI presents the studies that propose such custom evaluation metrics and lists the corresponding metrics introduced in each work. Studies that do not propose any custom evaluation metrics are omitted from this table.
V-C3 Libraries / Tools for Measurement
@@ V-C4 Hardware Configuration
Evaluation Hardware Category Primary Studies
On-premise (local machine/server) [105], [122], [21], [26], [124], [62], [111], [9]
@@ V-C5 Experimental Configuration
V-C6 Open-Source Availability
@@
Category Primary Studies
Open-source available [92], [105], [32], [21], [26], [24], [101], [62], [93], [111], [28], [9]
@@
V-C7 Benchmark Comparison and Usage
We categorize the evaluation data based on the type of benchmark or dataset used and the programming language considered. Table XVII summarizes the benchmarks and datasets, along with programming language information when explicitly reported. Programming languages are reported only when explicitly specified by the authors or when the language choice is central to the evaluation; for standardized benchmarks, the programming language is implicit and therefore omitted. Most studies primarily focus on Python, reflecting its widespread use in code generation research. Additionally, not all studies use standardized benchmarks, with several relying on custom data sources, indicating variability in evaluation practices.
@@ V-C7 Benchmark Comparison and Usage
V-D Model Adaptation
V-D1 Fine-Tuning
We categorize the studies based on the use of fine-tuning, i.e., whether model parameters are adapted to generate more sustainable code. Notably, none of the reviewed studies employ fine-tuning specifically for improving sustainability. This highlights an open research direction in leveraging fine-tuning to guide LLMs toward more energy-efficient and resource-aware code generation.
V-D2 Prompt Engineering
Table XVIII summarizes the use of prompt engineering techniques in the selected studies. A study is marked as “Yes” if it explicitly applies methods such as zero-shot, few-shot, chain-of-thought, or prompt variations, and as “No” if default prompts are used. Several studies employ prompt-based strategies, while others evaluate models without specific prompt engineering.
VI Results and Synthesis
VI-A RQ1: How effectively do LLMs produce code that aligns with sustainability principles in practice?
The purpose of this RQ is to evaluate whether LLMs can generate sustainable code in practice by examining the consistency of efficiency improvements across tasks and models, as well as key trade-offs such as energy efficiency and functional correctness. Although sustainability is often an explicit objective in the reviewed studies, the empirical evidence does not consistently show that LLM-generated code outperforms human baselines. Among the selected studies, 7 studies reported that LLM-generated code does not outperform human baselines in sustainability-oriented comparisons ( [124],[122],[24],[101],[62],[28],[9]). Within this group, 2 studies explicitly concluded that human-written code performs better than LLM-generated code ( [124, 28]). In addition, 3 studies characterized the outcome as inconsistent, indicating that any observed improvement depends on the task, or benchmark ([21, 125, 10]). The remaining 8 studies did not evaluate a direct human-versus-LLM comparison for sustainability outcomes ([92],[105], [32],[26],[97],[93],[111],[11].
Beyond the human-versus-LLM comparison, the reported sustainability outcomes are frequently context-dependent. Specifically, 14 studies indicate that efficiency gains are task-dependent, varying across benchmarks, problem types, or code characteristics ([124], [92], [122], [32], [21], [24], [101], [62], [97], [10], [93], [111] , [28], [11] ). Moreover, 11 studies report model-dependent outcomes, suggesting that sustainability performance can differ across LLMs and code assistants even under similar evaluation settings ([124], [21], [26], [101], [62], [10], [93], [11] , [9]) and one additional study reporting model sensitivity. Trade-offs are also a dominant theme in this literature with 16 studies explicitly discuss trade-offs ([124], [92], [105], [122], [32], [26], [24], [101], [125], [62], [97], [10], [93], [111] , [28], [9]). Within this group, 14 studies explicitly analyze the energy and accuracy trade-off ([124], [92], [105], [122], [32], [24], [101], [125], [62], [97], [10], [93], [111] , [9]), indicating that many studies treat sustainability as a multi-objective problem rather than a single-metric improvement. The remaining two studies discuss trade-offs but do not directly frame them as energy-versus-correctness ([26], [28]). Only two studies do not report an explicit trade-off discussion ([21], [11]). Overall, this pattern suggests that sustainability gains are frequently evaluated alongside potential costs such as correctness, and the results depend on how studies balance energy savings and correctness.
@@ VI-A RQ1: How effectively do LLMs produce code that aligns with sustainability principles in practice?
No [122], [24], [101], [125], [124], [62], [10], [111]
TABLE XVIII: Use of prompt engineering techniques
Most studies evaluate sustainability using a single final output per prompt/run, rather than generating multiple candidate solutions. Specifically, 12 studies report single-solution evaluation ([124], [92], [105], [122], [21], [24], [101], [125], [10], [111], [28], [11]). Only two studies employ Pass@k multi-sampling ([32], [62]), and among these, only one study explicitly reports an energy@k perspective ([32]). In addition, two studies generate multiple candidates through re-prompting without reporting Pass@k or energy@k metrics ([97], [9]), while one study does not clearly specify whether multi-sampling is used ([26]).
The energy overhead of generating multiple candidates is also largely under-reported. Only two studies explicitly state that producing additional candidates increases energy consumption ([32], [24]), and one study notes this effect qualitatively by indicating that additional refinement steps may increase overall cost, such as energy, latency ([92]). The remaining 15 studies do not quantify or report the energy impact of multi-candidate generation ([124], [105], [122], [21], [26], [101], [125], [62], [97], [10], [93], [111], [28], [11], [9]). Overall, these distributions indicate that, although multi-sampling and iterative prompting are common in practice, most studies assess sustainability using single-output settings and do not measure the additional energy overhead introduced by candidate sampling, limiting conclusions about end-to-end sustainability under realistic usage patterns.
Furthermore, the relationship between sustainability and correctness is mixed across the selected studies. Specifically, 8 studies report that energy-efficient outputs are also correct ([92], [105], [32], [62], [97], [93], [111], [9]), whereas 6 studies report that energy-efficient outputs are not consistently correct ([122], [24], [101], [125], [28], [11]); 2 studies report mixed outcomes ([124], [10]), and 2 studies do not evaluate this aspect ([21], [26]). Evidence regarding model-size effects remains limited. Only 2 studies report that larger models generate more energy-efficient code ([122], [10]), whereas 4 studies report contrary findings ([92], [101], [62], [11] ) and the remaining 13 studies do not evaluate model-size effects.
Notably, none of the studies discuss energy-per-token, indicating a consistent reporting gap. Regarding small models, 2 studies report that small models achieve roughly the same efficiency as large models ([26], [10]). In contrast, 5 studies indicate that small models do not achieve comparable efficiency ([124], [92], [122], [101], [62]). Moreover, 2 studies report inconsistent findings, where the outcome depends on the task or evaluation setting ([21], [11] ). The remaining 9 studies do not evaluate this aspect ([105], [32], [24], [125], [97], [93], [111] , [28], [9]). Finally, only 7 studies explicitly compare small and large models ([122], [21], [26], [62], [10], [93], [9]), whereas the remaining 11 studies do not evaluate this comparison ([124], [92], [105], [32], [24], [101], [125], [97], [111] , [28], [11] ). This limited coverage indicates that conclusions about sustainability effects across model scales remain constrained by the relatively small number of comparative studies.
VI-B RQ2: What metrics or parameters are used to evaluate sustainability?
This research question investigates how sustainability is evaluated in the literature by identifying the most commonly used metrics such as energy consumption, execution time, memory usage, carbon footprint, etc. Across the selected studies, sustainability evaluation is dominated by a small set of quantitative metrics. In all the studies, the sustainability of the generated code is measured using the metric energy consumption in 16 studies. Other aspects of sustainability are evaluated using the metric power consumption in two studies and the carbon footprint in one study. Two studies used custom evaluation metrics to measure the energy of the generated code, i.e., EnergyRatio, and Energy-Reduction@k. These studies are shown in Table. XII. In addition to energy consumption, 9 studies also reported memory-related metrics to account for the sustainability of the generated code. This suggests that sustainability has been measured using different evaluation metrics, however some metrics are more commonly used, such as energy consumption, compared to carbon footprint.
Finally, 4 studies introduce custom metrics ([124], [122], [32], [125]), typically motivated by capturing different dimensions of sustainability. RuntimeRatio and EnergyRatio compare the runtime and energy consumption of model-generated code against canonical human-written solutions, providing a relative efficiency measure. Energy-reduction@k estimates the expected energy savings when generating multiple candidate solutions and selecting the most energy-efficient valid one, while accounting for both performance improvement and correctness probability. Similarly, Green Capacity measures the overall sustainability improvement of optimized code across multiple performance dimensions, filtering out invalid or degraded solutions. Embodied Energy and Embodied Carbon capture the environmental impact associated with code generation or development infrastructure, whereas Operational Energy and Operational Carbon measure emissions during actual code execution. Carbon Intensity (CI) quantifies the amount of carbon emissions per unit of electricity consumed.
VI-C RQ3: What benchmarks, datasets, and tools are utilized to assess the sustainability of LLM-generated code?
The analysis of the selected studies shows that researchers rely on a diverse set of benchmarks and datasets to evaluate LLM-generated code, although several common sources emerge. A significant portion of the studies use programming challenge platforms, such as LeetCode, Codeforces, and HackerRank, where authors typically collect or design custom subsets of problems or code solutions from these websites rather than relying on a fixed benchmark dataset [122, 21, 24, 124, 10, 28, 11]. These platforms provide algorithmic programming tasks that are widely used to assess the correctness and efficiency of generated code. In addition to these sources, several studies employ established code generation benchmarks, including CodeXGLUE, HumanEval, EvoEval, and EffiBench, which are specifically designed for evaluating LLM-based code generation and software engineering tasks [105, 62, 111, 9]. Other works use large-scale code corpora and repositories, such as GitHub repositories, to analyze real-world code [101]. Some studies further rely on domain-specific benchmark suites, including HeCBench, XSBench, miniMDock, SciMark2, and DaCapoBench, which are commonly used in scientific computing and performance evaluation contexts [32, 93]. A few studies also introduce custom datasets or applications, such as the Energy-Language dataset or a CLI-based Connect-4 game [92, 26], while some studies do not clearly specify the dataset used [125, 97].
With respect to programming languages, the experiments predominantly focus on Python, followed by C++ and Java, with occasional inclusion of JavaScript and MATLAB. Several programming languages, such as Go and Solidity, are not represented in the reviewed studies and therefore remain unexplored in the context of sustainable code generation. Overall, the results indicate that most studies rely on custom datasets derived from programming challenge platforms or other problem sources, rather than consistently using standardized benchmarks for evaluating the sustainability of generated code.
The reviewed studies employ a variety of tools and libraries to measure the energy consumption of LLM-generated code, with most relying on software-based energy estimation frameworks. Commonly used tools include Intel RAPL, CodeCarbon, PyJoules, Perf, and PyNVML, which estimate energy consumption by accessing CPU or GPU performance counters and system-level statistics [92, 105, 122, 21, 24, 124, 111]. Additional system monitoring utilities such as Nvidia-smi, Rocm-smi, and Turbostat are also used to obtain hardware utilization and power-related metrics [26, 93]. However, very limited number of studies perform direct hardware-level energy measurements using external power monitoring devices. In particular, [28, 9] employ a Monsoon Power Monitor, specifically on Raspberry Pi platforms.
VI-D RQ4: How do prompting strategies or fine-tuning methods influence the sustainability of generated code?
Among the 18 primary studies, prompt engineering techniques explicitly targeting energy efficiency or sustainability are investigated in 10 studies ([92], [105], [32], [21], [26], [97], [93], [28], [11], [9]). The remaining 8 studies do not vary prompts for sustainability objectives ([122], [24], [101], [125], [124], [62], [10], [111]). Prompting strategies range from basic methods like zero-shot (only task description) and few-shot (with examples), to more structured approaches such as role-based prompting (assigning the model a specific role) and energy-specialised prompting (explicitly asking for efficient or low-resource outputs). More advanced techniques, including Chain-of-Thought (CoT) reasoning and iterative feedback loops, are also used to improve both performance and energy efficiency. Together, these prompting methods form a progression from simple instructions to more controlled and optimisation-focused designs.
Zero-shot prompting does not consistently reduce the energy consumption of generated code. Some studies report results similar to or even worse than baseline human-written or unoptimized code [124, 122]. In contrast, Chain-of-Thought (CoT) prompting shows more consistent improvements, particularly for small and medium-sized models. For example, [11] demonstrates that CoT outperforms baseline approaches for models such as Qwen2.5-Coder-3B and StableCode-3B, although it provides limited benefits for larger models like CodeLlama-7B and Phi-3-Mini-4K. Similarly, [9] reports energy savings with CoT, though the results vary across hardware platforms.
Explicit energy-aware prompting produces mixed results across studies. Some works show measurable reductions in energy consumption [101, 97, 9], with one study reporting up to a 36% reduction when combined with iterative feedback [97]. However, other studies find that such prompts can lead to neutral or inconsistent outcomes compared to baseline prompting [122, 21]. Iterative prompting strategies, which incorporate self-correcting feedback loops and LLM-as-a-judge mechanisms, show the strongest and most consistent improvements in the studies that adopt them. These approaches achieve significant reductions in energy consumption while maintaining functional correctness, although they require multiple generation steps [32, 97].
Overall, no single prompting strategy consistently outperforms others across all settings. The effectiveness of a given approach depends on factors such as the model, the nature of the task, and the execution environment. In general, more structured prompting techniques, such as Chain-of-Thought, tend to perform better than simpler approaches like zero-shot in certain scenarios, while specialized and iterative prompting methods often provide additional benefits in practical, real-world applications.
VII Future Research Opportunities
@@
VII-A Expanding Sustainability Beyond Code Generation
@@
VII-B Underexplored Application Domains
@@
Future research should explore sustainability-aware code generation across a broader range of application domains. In particular, mobile and client-side applications, low-level and systems software, web and cloud infrastructure, and IoT and embedded systems are not adequately represented in the selected studies. These domains are especially important from a sustainability perspective, as energy efficiency directly affects battery life, system performance, scalability, and overall environmental impact. For example, embedded and IoT systems operate under strict energy constraints, low-level software can significantly influence hardware efficiency, and cloud-based services contribute substantially to large-scale energy consumption. The limited attention to these areas highlights the need for future studies to evaluate LLM-generated code in more realistic, and deployment-oriented environments.
VII-C Expanding Sustainability Evaluation Across Programming Languages
@@
This direction is particularly important from sustainability perspective, as energy efficiency can vary significantly across programming languages and execution environments. Different languages introduce different abstractions, memory management mechanisms, and runtime behaviors, all of which influence energy consumption. Moreover, language choice is often tightly coupled with the target platform, such as mobile devices, cloud infrastructure, edge systems, blockchain networks, or networked distributed systems, which further shapes optimization opportunities, runtime characteristics, and energy profiles.
@@
VII-D Toward Sustainability-Aware LLM Optimization
Another important direction for future research is the development of sustainability-aware fine-tuning and model adaptation techniques. Our review shows that no study used model adaptation techniques, e.g., fine-tuning with an explicit focus on improving energy efficiency. This suggests that most existing work measures sustainability at the output level without modifying the underlying model to generate more energy-efficient code.
Future research should explore energy-aware fine-tuning strategies specifically tailored for sustainable code generation and optimization tasks. Instead of only measuring the energy consumption of generated code after inference, models should be trained to directly produce energy-efficient implementations. For example, training objectives can be extended to penalize inefficient code patterns, high-complexity logic, or unnecessary computational operations, encouraging the generation of optimized and resource-efficient solutions. Green reinforcement learning approaches may also be applied, where models receive reward signals based on improvements in runtime efficiency, reduced energy consumption, or lower carbon footprint of the generated code. Moreover, parameter-efficient adaptation methods can further reduce training and deployment cost, making sustainability-aware code generation more practical. By embedding sustainability objectives directly into model training and inference processes, future research can enable LLMs to proactively generate energy-efficient code rather than relying solely on post-hoc energy evaluation.
VII-E Limited Focus on Small Models
@@
VII-F Need for Sustainability-Oriented Benchmarks
Another important observation from this study is that none of the selected works introduce a benchmark specifically designed to evaluate sustainability. Instead, all studies use existing general-purpose benchmarks and add sustainability-related metrics, such as energy consumption or carbon footprint, as additional evaluation measures. This indicates a clear research gap in the field. Future research should focus on developing standardized benchmarks dedicated to sustainability-aware code generation. These benchmarks should include tasks and evaluation methods that measure energy efficiency, carbon impact, resource usage, and the ability of models to produce energy-efficient solutions. Creating such benchmarks would improve comparability across studies and support more rigorous evaluation of sustainable code generation approaches.
VII-G Energy Measurement and Hardware Evaluation Gaps